# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 8 Impact of the Gaussian likelihood assumption
One of the main aims of this work is to determine whether the standard assumption of the likelihood of cosmic shear two-point statistics being Gaussian holds under realistic conditions while including systematics. To test this assumption, we perform the same analysis of a mock data vector using SBI described in Sect. 5 and Appendix E again with different inference methods.
As shown in Fig. 17, we compare the posterior from the SBI analysis learnt from forward simulations based on the anisotropic systematics model to a standard Gaussian likelihood analysis and an SBI analysis which forces the learnt likelihood to be Gaussian. The standard analysis assumes a Gaussian likelihood characterised by the same numerical covariance from 1,600 realisations of the anisotropic systematics model fixed at the fiducial cosmology in Table 2. The model evaluations are calculated from analytical theory for the cosmic shear signal as described in Appendix D, and the sampling is done with the nautilus sampler (Lange, 2023) over the same 12-dimensional prior space defined in Table 2. In contrast, the Gaussian SBI analysis makes the exact same modelling choices as the main SBI analysis described Sect. 5 and Appendix E with one change: instead of using ensembles of MAFs made up of MADEs as the architecture of the neural density estimators when learning the effective likelihood, we use an ensemble of three single-component Mixture Density Networks (MDNs; Bishop 1994) to learn the likelihood from the 18,000 realisation of the anisotropic systematics model. As a consequence, the ensemble of MDNs forces the learnt likelihood to be described by a single multi-variate Gaussian distribution. Notably, this is different from a standard Gaussian likelihood analysis with a Gaussian likelihood in that the Gaussian SBI analysis allows the covariance of the likelihood to vary with cosmology.
@@
At the same time, we find that the standard Gaussian likelhood analysis has a noticeably different posterior to both SBI analyses despite the underlying data vector being identical, and the covariance being sampled from the same anisotropic systematics model. In particular, the $S_{8}$ marginal is appreciably narrower in the Gaussian likelihood analysis than in the SBI analyses. In the noiseless case shown in Fig. 17, we find that the SBI gives $S_{8}=0.743^{+0.034}_{-0.031}$, while the standard Gaussian likelihood analysis gives $S_{8}=0.733^{+0.034}_{-0.024}$, i.e. the $1\sigma$ confidence interval in $S_{8}$ is ${\sim}11\%$ narrower than in the SBI analysis. When adding realistic noise to the data vector from the same seed, we find that the SBI recovers $S_{8}=0.762^{+0.032}_{-0.027}$, while the Gaussian likelihood analysis recovers $S_{8}=0.744^{+0.031}_{-0.022}$, i.e. the $1\sigma$ confidence interval in $S_{8}$ is also ${\sim}10\%$ narrower than in the SBI analysis. Additionally, when noise is introduced, the Gaussian likelihood analysis appears more biased with respect to the truth ($S_{8}=0.756$) than the SBI.
@@
To investigate the origin of this effect further, we sample the learnt likelihood from the full SBI analysis directly at different values of $S_{8}$ in Fig. 18. In this plot, we show the seven-dimensional score-compressed data space and label each element of the data vector according to the model parameter with which they are most correlated. Then, we vary the value of $S_{8}$ while keeping all other parameters fixed at the fiducial cosmology, and plot the likelihood of the data given the chosen model parameters. We find that as $S_{8}$ increases the marginal likelihoods of the elements of the score-compressed data vector increase linearly. In particular, the data elements most correlated with $\omega_{c}$, $\omega_{b}$, $n_{s}$, and $h_{0}$ show the strongest linear correlation with $S_{8}$. The implication of this is that the anisotropic systematics model entails that data at higher assumed $S_{8}$ values has larger uncertainties than data modelled for low $S_{8}$ values, and therefore the learnt likelihood is different from a Gaussian likelihood with the covariance evaluated at a fixed cosmology. This effect appears to drive the increase in the uncertainty on $S_{8}$ of the SBI-based posteriors with respect to the Gaussian likelihood analysis in Fig. 17.
@@
In Appendix F, we conduct a detailed investigation into the physical origin of this observed cosmology dependence in the learnt likelihood of cosmic shear two-point statistics. We find that, as has been noted in previous work (Eifler et al., 2009; Reischke et al., 2017), the cosmic variance of the cosmic shear two-point signal (at least, for $76\leq\ell\leq 1500$) is measurably cosmology-dependent, and mostly $S_{8}$-dependent. We also determine that this effect is expected to be important enough to become detectable in most tomographic bin combinations of the KiDS-1000 gold sample in the regime of $\ell\sim 10^{2}$. Since the cosmic variance is imprinted into the forward simulations at the step of sampling log-normal random matter fields (see Sect. 3.4), the cosmology dependence of the cosmic variance is present in any forward model presented in this work irrespective of the observational systematic effects considered (i.e. both the anisotropic systematics and the standard isotropic systematics model are affected by this).
@@
With this being the case, the noise of the measured cosmic shear pseudo-Cls may contain some information which may impact the inferred cosmology from the SBI. Hence, it is important to validate whether the variance realised by log-normal random simulations does not bias the cosmological inference. Firstly, it has been found in previous studies that log-normal simulations can incorporate accurate higher-order statistics into the sampled matter, galaxy and shear fields when compared to N-body dark matter simulations (Hall & Taylor, 2022; Piras et al., 2023), so log-normal fields can incorporate three- or four-point statistics to model the variance of two-point statistics. Additionally, we find in our own testing that a sample covariance of cosmic shear two-point statistics in KiDS-1000 as generated from log-normal simulations agrees within $10\%$ with analytical covariance matrices computed from analytical considerations (Joachimi et al. 2021, Reischke et al. in prep.), so the variance of the cosmic shear signal as modelled within KiDS-SBI is consistent with analytical theory.
@@ 9.2 Cosmological parameter constraints
Figure 19: Posterior contours of the main constrained cosmological parameters from the KiDS-SBI analysis of the KiDS-1000 cosmic shear data assuming the anisotropic systematics model (in blue), which incorporates additional systematics such as variable depth and shear biases, compared against posterior contours from other analyses. In pink, we show the contours for the equivalent analysis assuming a Gaussian likelihood sampled with nautilus (Lange, 2023). In orange, we show the posterior from the same data while assuming the standard isotropic systematics model, which considers the systematic effects which are typically modelled in standard cosmic shear analyses. In black, we show the posterior sampled with nautilus based on a Gaussian likelihood defined by the standard isotropic systematics model. The green contour shows the posterior from the cosmic microwave background constraints from the TT,TE,EE+lowE modes (Planck Collaboration et al., 2020). Note that the Planck contours do not have any marginals in $A_{\mathrm{IA}}$ as the CMB is not sensitive to the intrinsic alingments of galaxies.
@@
As also shown in Fig. 19, we find that when re-analysing the KiDS-1000 data assuming a Gaussian likelihood instead, we find results consistent with the testing conducted on mocks (see Fig. 17): the Gaussian likelihood assumption causes $S_{8}$ to be overconstrained by ${\sim}10\%$ with respect to the SBI irrespective of how the observational systematics are modelled as it neglects to consider the cosmology dependence of cosmic variance. Additionally, the posterior mean marginal in $S_{8}$ assuming the Gaussian likelihood shifts from $0.764^{+0.031}_{-0.025}$ when deriving the sample covariance from the standard isotropic systematics model to a value of $0.725^{+0.034}_{-0.023}$ when assuming the anisotropic systematics model in the covariance instead. Note that both assume the same modelling in the signal described in Appendix D.
We highlight that the Gaussian likelihood analysis based on the standard isotropic systematics model agrees well with the constraints from Asgari et al. (2021) where $S_{8}=0.758^{+0.017}_{-0.026}$. We note that the constraints in our work are broader despite making consistent modelling assumptions. This is caused by the use of a different MCMC sampler. Here we have employed the nautilus sampler (Lange, 2023), while in Asgari et al. (2021) the multinest sampler (Feroz et al., 2009) was used. It has been found that multinest can be overconfident by approximately ${\sim}10\%$ (Lemos et al., 2023b) which we also find in our analysis and show in Appendix H. When accounting for the difference in sampler used, we find that the Gaussian likelihood constraints from the isotropic systematics model agree exceptionally well with the results from Asgari et al. (2021).
@@ 10 Conclusions
We presented a new suite of stochastic forward simulations of cosmic shear observables within KiDS-SBI which take into consideration all the systematic effects which are typically modelled in a weak lensing analysis plus additional systematics which have the potential to significantly bias the signal. At the same time, KiDS-SBI introduces considerable improvements in accuracy and performance with respect to similar simulation environments in previous work (Joachimi et al., 2021): non-Limber projections of angular power spectra (with the new Levin module), higher resolution along the line-of-sight, inclusion of field-level intrinsic alignments, consideration of variable depth and anisotropic shear biases, and significant computational performance improvements.
@@
On the basis of these forward simulations, we implement the first simulation-based inference analysis of large-scale structure observables as measured by KiDS-1000 with the same level of complexity as other standard analyses in the field. Thus, this enables us to fully propagate all uncertainties from the data vector all the way to the inferred cosmological parameters in a Bayesian way. The analysis is based on neural density estimators which learn the effective likelihood through the mapping between the model parameters and the forward-simulated data vectors. We conduct inference of the cosmological parameters in flat $\Lambda$CDM from cosmic shear two-point statistics, specifically, pseudo-Cls. We assume a forward model consistent with previous KiDS-1000 cosmic shear analyses (Asgari et al., 2021; Loureiro et al., 2022) which additionally includes variable depth and anisotropic shear biases: the anisotropic systematics model. Variable depth accounts for the anisotropy in the galaxy selection function in KiDS-1000 and how it impacts the measured galaxy densities, shapes and redshifts. The inclusion of anisotropic shear biases models the angular variation of the residual bias in the galaxy shapes due to variations in the point-spread function of the telescope from the mean model used in the shape calibration.
@@
To assess the quality of any best fit obtained from KiDS-SBI, we also introduce a Bayesian goodness-of-fit measure based on Gelman et al. (1996). From a small number of realisations of the forward model, we characterise a distribution of $\chi^{2}$ values for the given model. By determining where the $\chi^{2}$ of the best fit on the KiDS-1000 data lies, we evaluate the quality of the fit, and whether the data is a feasible realisation of the forward model. From this, we conclude that the KiDS-1000 data is fitted well by the anisotropic systematics model at the maximum a posteriori cosmology (with a probability-to-exceed of 0.42).
@@
Through the testing conducted in this work and in a previous paper (Lin et al., 2023), we determine that KiDS-SBI recovers accurate and robust posteriors of the 12-dimensional space of cosmological and astrophysical parameters. At the same time, we find that the likelihood learnt from KiDS-SBI forward simulations is well approximated by a Gaussian distribution at a fixed cosmology. However, we determine that the uncertainty on the cosmic shear signal measurably varies with the assumed cosmological parameters which are driven by the cosmic variance scaling with the underlying value of $S_{8}$ as has been noted in previous work (Eifler et al., 2009). Consequently, we find that the uncertainty of the two-point statistics in KiDS-1000 is measurably cosmology-dependent which increases the nominal uncertainty of $S_{8}$ by ${\sim}10\%$. Hence, the standard assumption of the covariance of a Gaussian likelihood being fixed for all parameter points in the prior space does not necessarily hold. Therefore, such an assumption may not be valid for current Stage-III galaxy surveys, and it may be also be relevant for upcoming Stage-IV galaxy surveys, such as Euclid (Laureijs et al., 2011) or Rubin (LSST Science Collaboration et al., 2009), which will further improve the precision of cosmic shear measurements.
@@
To evaluate and isolate the impact of the inclusion of variable depth and shear biases in the anisotropic systematics model, we reproduce the KiDS-SBI analysis with a second model which makes the same assumptions, but does not consider these specific systematic effects: the standard isotropic systematics model (which is more in line with previous KiDS-1000 analyses, e.g. Asgari et al. 2021). We find that the inclusion of both systematics in the forward model increases the $S_{8}$ marginal by $0.9\sigma$ or $5\%$ and decreases the marginal $1\sigma$ confidence intervals by $6\%$. This reflects the fact that variable depth and PSF shear variation bias can significantly alter the uncertainty modelling despite having a negligible impact on the shear signal at the precision levels of KiDS-1000. The KiDS-1000 data appears to be consistent with the presence of these effects since the anisotropic systematics model provides a good fit to the data, but when examining the Bayesian evidence, we do not find a clear preference between the two models by the data. Nevertheless, both systematic effects have been independently measured in the KiDS-1000 weak lensing data (Joachimi et al., 2021; Giblin et al., 2021).
@@
This work stresses the importance of testing the common assumption of Gaussian likelihoods in realistic conditions. At the same time, we also highlight the potential of simulation-based inference to rigorously perform Bayesian uncertainty propagation even when the models of the uncertainty of the signal or the systematics are analytically intractable. SBI with likelihood-based neural density estimation comes with the added advantage that it does not require the expensive re-evaluation of the posterior distribution if the input measurements or priors are updated, as would be the case for a standard MCMC analysis. Since the likelihood for a given forward model is learnt independently of the input data and the priors, SBI can significantly improve the efficiency of inference analysis where the forward model is not changed between measurements. All these advantages showcase the potential of simulation-based inference becoming the method of choice in future parameter inference and model testing analyses in observational cosmology.
@@
We conclude that SBI is a powerful tool which may be able to tackle many of the physical and statistical challenges which future galaxy surveys will face, as their observations become more precise, less limited by statistical noise, and more limited by the accurate modelling of systematics. To address this need, future work should expand upon the forward model in KiDS-SBI to also include other probes of large-scale structure such as galaxy clustering and galaxy-galaxy lensing, while incorporating the modelling for the relevant systematic effects, such as magnification bias, non-linear galaxy bias and field-level baryonic feedback.

### v2

@@ 8 Impact of the Gaussian likelihood assumption
One of the main aims of this work is to determine whether the standard assumption of the likelihood of cosmic shear two-point statistics being Gaussian holds under realistic conditions while including systematics. To test this assumption, we performed the same analysis of a mock data vector using SBI (as described in Sect. 5 and Appendix E) again, but with different inference methods.
As shown in Fig. 16, we compare the posterior from the SBI analysis learnt from forward simulations based on the anisotropic systematics model to a standard Gaussian likelihood analysis and an SBI analysis which forces the learnt likelihood to be Gaussian. The standard analysis assumes a Gaussian likelihood characterised by the same numerical covariance from 1,600 realisations of the anisotropic systematics model fixed at the fiducial cosmology in Table 18. The model evaluations are calculated from analytical theory for the cosmic shear signal as described in Appendix D, and the sampling is done with the nautilus sampler (Lange, 2023) over the same 12D prior space defined in Table 18. In contrast, the Gaussian SBI analysis makes the exact same modelling choices as the main SBI analysis described Sect. 5 and Appendix E with one change: instead of using ensembles of MAFs made up of MADEs as the architecture of the neural density estimators when learning the effective likelihood, we use an ensemble of three single-component Mixture Density Networks (MDNs; Bishop 1994) to learn the likelihood from the 18,000 realisation of the anisotropic systematics model. As a consequence, the ensemble of MDNs forces the learnt likelihood to be described by a single multi-variate Gaussian distribution. Notably, this is different from a standard Gaussian likelihood analysis with a Gaussian likelihood in that the Gaussian SBI analysis allows the covariance of the likelihood to vary with cosmology.
@@
At the same time, we find that the standard Gaussian likelihood analysis has a noticeably different posterior to both SBI analyses despite the underlying data vector being identical, and the covariance being sampled from the same anisotropic systematics model. In particular, the $S_{8}$ marginal is appreciably narrower in the Gaussian likelihood analysis than in the SBI analyses. In the noiseless case shown in Fig. 16, we find that the SBI gives $S_{8}=0.743^{+0.034}_{-0.031}$, while the standard Gaussian likelihood analysis gives $S_{8}=0.733^{+0.034}_{-0.024}$, specifically the $1\sigma$ confidence interval in $S_{8}$ is ${\sim}11\%$ narrower than in the SBI analysis. When adding realistic noise to the data vector from the same seed, we find that the SBI recovers $S_{8}=0.762^{+0.032}_{-0.027}$, while the Gaussian likelihood analysis recovers $S_{8}=0.744^{+0.031}_{-0.022}$, specifically the $1\sigma$ confidence interval in $S_{8}$ is also ${\sim}10\%$ narrower than in the SBI analysis. Additionally, when noise is introduced, the Gaussian likelihood analysis appears more biased with respect to the truth ($S_{8}=0.756$) than the SBI.
@@
To investigate the origin of this effect further, we sample the learnt likelihood from the full SBI analysis directly at different values of $S_{8}$ in Fig. 17. In this plot, we show the 7D score-compressed data space and label each element of the data vector according to the model parameter with which they are most correlated. Then, we vary the value of $S_{8}$ while keeping all other parameters fixed at the fiducial cosmology, and plot the likelihood of the data given the chosen model parameters. We find that as $S_{8}$ increases the marginal likelihoods of the elements of the score-compressed data vector increase linearly. In particular, the data elements most correlated with $\omega_{c}$, $\omega_{b}$, $n_{s}$, and $h_{0}$ show the strongest linear correlation with $S_{8}$. The implication of this is that the anisotropic systematics model entails that data at higher assumed $S_{8}$ values has larger uncertainties than data modelled for low $S_{8}$ values, and therefore the learnt likelihood is different from a Gaussian likelihood with the covariance evaluated at a fixed cosmology. This effect appears to drive the increase in the uncertainty on $S_{8}$ of the SBI-based posteriors with respect to the Gaussian likelihood analysis in Fig. 16.
@@
In Appendix F, we conduct a detailed investigation into the physical origin of this observed cosmology dependence in the learnt likelihood of cosmic shear two-point statistics. We find that, as has been noted in previous work (Eifler et al., 2009; Reischke et al., 2017), the cosmic variance of the cosmic shear two-point signal (at least, for $76\leq\ell\leq 1500$) is measurably cosmology-dependent, and mostly $S_{8}$-dependent. We also determine that this effect is expected to be important enough to become detectable in most tomographic bin combinations of the KiDS-1000 gold sample in the regime of $\ell\sim 10^{2}$. Since the cosmic variance is imprinted into the forward simulations at the step of sampling log-normal random matter fields (see Sect. 3.4), the cosmology dependence of the cosmic variance is present in any forward model presented in this work irrespective of the observational systematic effects considered (namely both the anisotropic systematics and the standard isotropic systematics model are affected by this).
@@
With this being the case, the noise of the measured cosmic shear pseudo-Cls may contain some information which may impact the inferred cosmology from the SBI. Hence, it is important to validate whether the variance realised by log-normal random simulations does not bias the cosmological inference. Firstly, it has been found in previous studies that log-normal simulations can incorporate accurate higher-order statistics into the sampled matter, galaxy and shear fields when compared to N-body dark matter simulations (Hall & Taylor, 2022; Piras et al., 2023), so log-normal fields can incorporate three- or four-point statistics to model the variance of two-point statistics. Additionally, we find in our own testing that a sample covariance of cosmic shear two-point statistics in KiDS-1000 as generated from log-normal simulations agrees within $10\%$ with analytical covariance matrices computed from analytical considerations (Joachimi et al., 2021; Reischke et al., 2024), so the variance of the cosmic shear signal as modelled within KiDS-SBI is consistent with analytical theory.
@@ 9.2 Cosmological parameter constraints
Figure 18: Posterior contours of the main constrained cosmological parameters from the KiDS-SBI analysis of the KiDS-1000 cosmic shear data assuming the anisotropic systematics model (in blue), which incorporates additional systematics such as variable depth and shear biases, compared against posterior contours from other analyses. In pink, we show the contours for the equivalent analysis assuming a Gaussian likelihood sampled with nautilus (Lange, 2023). In orange, we show the posterior from the same data while assuming the standard isotropic systematics model, which considers the systematic effects which are typically modelled in standard cosmic shear analyses. In black, we show the posterior sampled with nautilus based on a Gaussian likelihood defined by the standard isotropic systematics model. The green contour shows the posterior from the cosmic microwave background constraints from the TT,TE,EE+lowE modes (Planck Collaboration et al., 2020). We note that the Planck contours do not have any marginals in $A_{\mathrm{IA}}$ as the CMB is not sensitive to the intrinsic alingments of galaxies.
Table 3: Main inferred cosmological parameters varied within the anisotropic systematics model from the KiDS-1000 gold sample.
@@
As also shown in Fig. 18, we find that when re-analysing the KiDS-1000 data assuming a Gaussian likelihood instead, we find results consistent with the testing conducted on mocks (see Fig. 16): the Gaussian likelihood assumption causes $S_{8}$ to be overconstrained by ${\sim}10\%$ with respect to the SBI irrespective of how the observational systematics are modelled as it neglects to consider the cosmology dependence of cosmic variance. Additionally, the posterior mean marginal in $S_{8}$ assuming the Gaussian likelihood shifts from $0.764^{+0.031}_{-0.025}$ when deriving the sample covariance from the standard isotropic systematics model to a value of $0.725^{+0.034}_{-0.023}$ when assuming the anisotropic systematics model in the covariance instead. We note that both assume the same modelling in the signal described in Appendix D.
We highlight that the Gaussian likelihood analysis based on the standard isotropic systematics model agrees well with the constraints from Asgari et al. (2021) where $S_{8}=0.758^{+0.017}_{-0.026}$. We note that the constraints in our work are broader despite making consistent modelling assumptions. This is caused by the use of a different MCMC sampler. Here we have employed the nautilus sampler (Lange, 2023), while in Asgari et al. (2021) the multinest sampler (Feroz et al., 2009) was used. It has been found that multinest can be overconfident by approximately ${\sim}10\%$ (Lemos et al., 2023b) which we also find in our analysis and show in Appendix G. When accounting for the difference in sampler used, we find that the Gaussian likelihood constraints from the standard isotropic systematics model agree exceptionally well with the results from Asgari et al. (2021).
@@ 10 Conclusions
We present a new suite of stochastic forward simulations of cosmic shear observables within KiDS-SBI, which take into consideration all the systematic effects which are typically modelled in a weak lensing analysis plus additional systematics that have the potential to significantly bias the signal. At the same time, KiDS-SBI introduces considerable improvements in accuracy and performance with respect to similar simulation environments in previous works (Joachimi et al., 2021): non-Limber projections of angular power spectra (with the new Levin module), higher resolution along the line of sight, inclusion of field-level intrinsic alignments, consideration of variable depth and anisotropic shear biases, and significant computational performance improvements.
@@
On the basis of these forward simulations, we implemented the first simulation-based inference analysis of large-scale structure observables as measured by KiDS-1000 with the same level of complexity as other standard analyses in the field. This has enabled us to fully propagate all uncertainties from the data vector all the way to the inferred cosmological parameters in a Bayesian way. The analysis is based on neural density estimators which learn the effective likelihood through the mapping between the model parameters and the forward-simulated data vectors. We conducted an inference of the cosmological parameters in flat $\Lambda$CDM from cosmic shear two-point statistics, specifically, pseudo-Cls. We assumed a forward model consistent with previous KiDS-1000 cosmic shear analyses (Asgari et al., 2021; Loureiro et al., 2022), which additionally includes variable depth and anisotropic shear biases: the anisotropic systematics model. Variable depth accounts for the anisotropy in the galaxy selection function in KiDS-1000 and how it impacts the measured galaxy densities, shapes, and redshifts. The inclusion of anisotropic shear biases models the angular variation of the residual bias in the galaxy shapes due to variations in the point-spread function of the telescope from the mean model used in the shape calibration.
@@
To assess the quality of any best fit obtained from KiDS-SBI, we also introduced a Bayesian goodness-of-fit measurement based on Gelman et al. (1996). From a small number of realisations of the forward model, we characterised a distribution of $\chi^{2}$ values for the given model. By determining where the $\chi^{2}$ of the best fit on the KiDS-1000 data lies, we evaluated the quality of the fit and whether the data stand as a feasible realisation of the forward model. On this basis, we were able to conclude that the KiDS-1000 data is fitted well by the anisotropic systematics model at the maximum a posteriori cosmology (with a probability-to-exceed of 0.42).
@@
Through the testing conducted in this work and in a previous paper (Lin et al., 2023), we determine that KiDS-SBI recovers accurate and robust posteriors of the 12D space of cosmological and astrophysical parameters. At the same time, we find that the likelihood learnt from KiDS-SBI forward simulations is well approximated by a Gaussian distribution at a fixed cosmology. However, we determined that the uncertainty on the cosmic shear signal measurably varies with the assumed cosmological parameters, which are driven by the cosmic variance scaling with the underlying value of $S_{8}$, as noted in previous works (Eifler et al., 2009). Consequently, we find that the uncertainty of the two-point statistics in KiDS-1000 is measurably cosmology-dependent which increases the nominal uncertainty of $S_{8}$ by ${\sim}10\%$. Hence, the standard assumption of the covariance of a Gaussian likelihood being fixed for all parameter points in the prior space does not necessarily hold. Therefore, such an assumption may not be valid for current Stage-III galaxy surveys and this issue may be also be relevant for upcoming stage-IV galaxy surveys, such as Euclid (Laureijs et al., 2011) or Rubin (LSST Science Collaboration et al., 2009). Addressing this assumption further will improve the precision of cosmic shear measurements.
@@
To evaluate and isolate the impact of the inclusion of variable depth and shear biases in the anisotropic systematics model, we reproduce the KiDS-SBI analysis with a second model which makes the same assumptions, but does not consider these specific systematic effects: the standard isotropic systematics model (which is more in line with previous KiDS-1000 analyses, such as Asgari et al. 2021). We find that the inclusion of both systematics in the forward model increases the $S_{8}$ marginal by $0.9\sigma$ or $5\%$ and decreases the marginal $1\sigma$ confidence intervals by $6\%$. We find that this is not driven by these systematics’ impact on the signal but, rather, on their impact on the uncertainty model. The variable depth alone appears to impact $S_{8}$ by approx. $0.38\sigma$ when comparing the Gaussian likelihood results from Loureiro et al. (2022) with the constraints from the standard isotropic systematics model. The variations in the PSF modelling residuals appear to account for approx. $0.36\sigma$ in the shift in $S_{8}$ when comparing the Gaussian likelihood results from from the anisotropic systematics model with the constraints from Loureiro et al. (2022). However, we note that the PSF variations and the variable depth are known to be correlated as both are influenced by some of the same observational systematics, such as atmospheric seeing or thermal effects in the telescope optics; thus, it may not be physically sound to consider these effects separately.
In any case, this reflects the fact that observational anisotropies, such as variable depth and PSF shear variation bias, can significantly alter the uncertainty modelling despite having a negligible impact on the shear signal at the precision levels of KiDS-1000. The KiDS-1000 data appears to be consistent with the presence of these effects since the anisotropic systematics model provides a good fit to the data, but when examining the Bayesian evidence, we do not find a clear preference between the two models by the data. Nevertheless, both systematic effects have been independently measured in the KiDS-1000 weak lensing data (Joachimi et al., 2021; Giblin et al., 2021).
@@
This work stresses the importance of testing the common assumption of Gaussian likelihoods in realistic conditions. At the same time, we also highlight the potential of simulation-based inference to rigorously perform Bayesian uncertainty propagation even when the models of the uncertainty of the signal or the systematics are analytically intractable. The approach using SBI with likelihood-based neural density estimation comes with the added advantage that it does not require the expensive re-evaluation of the posterior distribution if the input measurements or priors are updated, as would be the case for a standard MCMC analysis. Since the likelihood for a given forward model is learnt independently of the input data and the priors, SBI can significantly improve the efficiency of inference analysis where the forward model is not changed between measurements. All these advantages showcase the potential of simulation-based inference becoming the method of choice in future parameter inference and model testing analyses in observational cosmology.
@@
We conclude that SBI is a powerful tool that may be able to tackle many of the physical and statistical challenges faced by future galaxy surveys, as their observations become more precise, less limited by statistical noise, and more limited by the accurate modelling of systematics. To address this need, future works should expand upon the forward model in KiDS-SBI to also include other probes of large-scale structure such as galaxy clustering and galaxy-galaxy lensing, while incorporating the relevant modelling for the relevant systematic effects, such as magnification bias, non-linear galaxy bias, and field-level baryonic feedback.